Question 2

A spring-mass system, with a spring stiffness of 5,000 N/m, is subjected to a harmonic force of 

magnitude 30 N and frequency 20 Hz. The mass is found to vibrate with an amplitude of 0.2 m.

Assuming that vibration starts from rest (x0 = ẋ0 = 0), determine the mass of the system. [10]


1
Expert's answer
2021-10-04T15:30:30-0400

The static deflection

"δ_{st} = \frac{F_0}{k} \\

= \frac{30 \;N}{5000 \;M/m} \\

= 0.006 \;m"

The frequency of the motion

"ω= 2 \pi f \\

= 2 \pi (20) \\

= 125.66 \;rad/s"

The natural frequency of motion

"X = \frac{δ_{st}}{(1-(\frac{ω}{ω_n})^2)} \\

0.2 = \frac{0.006}{(1 -(\frac{125.66}{ω_n})^2)} \\

1 -(\frac{125.66}{ω_n})^2 = 0.03 \\

\frac{125.66}{ω_n} = 0.9848 \\

ω_n = 127.58 \;rad/s"

The mass of the system:

"ω_n = \sqrt{\frac{k}{m}} \\

127.58 = \sqrt{\frac{5000}{m}} \\

m = 0.3071 \;kg"